# Right Splitting, Galois Correspondence, Galois Representations and Inverse Galois Problem

### v1

@@ 1. Introduction
In this article, we establish some instances of Inverse Galois problem. In Sec. 2, through Galois correspondence and right splitting of some exact sequences of groups, we obtain the main results Thm. 2.4, Thm. 2.6. We then apply these general results to the case in [1] and obtain interesting consequences as corollaries.
In Sec. 3, using the algebraic operations on Galois representations and right splitting of some exact sequences of groups, we obtain Thm. 3.7, Thm. 3.8 and Thm. 3.19 as the main results. We then apply these general results to the case in [8] and obtain equally interesting consequences in Thm. 3.12 and the corollary of Thm. 3.19. Right Splitting of exact sequences of groups is the common thread that runs through the whole article.
@@ 2.1. Right splitting and group theoretic results
Proposition 2.1.
@@
where $\pi$ is the quotient map $G\rightarrow G/[G,G]$.
Proof.
Suppose $G/[G,G]$ is a cyclic group of order $m$ and $n\mid m$. Hence $G/[G,G]$ has a unique index-n cyclic subgroup, call it $H^{\prime}$, which is precisely given by $H^{\prime}=\{z^{n}\ |\ z\in G/[G,G]\}$, which is generated by $y^{n}$, where $y$ is a generator of $G/[G,G]$.
Since $\pi$ is surjective, we have one to one correspondence between subgroups of $G/[G,G]$ and subgroups of $G$ containing $[G,G]$. Let $H=\pi^{-1}(H^{\prime})=\{x\in G|\pi(x)$ is an $n$-th power in $G/[G,G]\}$. Hence it is the unique index-n normal subgroup of $G$ containing $[G,G]$.
Let $K$ be any index-$n$ normal subgroup of $G$ such that the quotient is abelian. Consider the quotient map $\rho:G\rightarrow G/K$. Since ${\rm Image}(\rho)$ is abelian, hence $[G,G]\subset Ker(\rho)=K$. Hence $K=H$.
Also we have that $G/H$ is isomorphic to $(G/[G,G])/H^{\prime}$.  Hence it is cyclic.
Conversely, suppose $G/[G,G]$ is not cyclic. Hence, by structure theorem for abelian groups,
$G/[G,G]\cong\mathbb{Z}/a_{1}\mathbb{Z}\times\mathbb{Z}/a_{2}\mathbb{Z}\times G^{\prime}$
for some group $G^{\prime}$, $a_{1}\mid a_{2}$ and $a_{1}>1$. Since $\mathbb{Z}/a_{2}\mathbb{Z}$ is cyclic and $a_{1}\mid a_{2}$, we get a unique index-$a_{1}$ subgroup $G^{\prime\prime}$ of $\mathbb{Z}/a_{2}\mathbb{Z}$. Hence $\mathbb{Z}/a_{2}\mathbb{Z}\times G^{\prime}$ and $\mathbb{Z}/a_{1}\mathbb{Z}\times G^{\prime\prime}\times G^{\prime}$ are distinct index-$a_{1}$ subgroups of $G/[G,G]$. Hence we don’t get a unique index-$a_{1}$ subgroup $H^{\prime}$ of $G/[G,G]$. By the one to one correspondence between subgroups of $G/[G,G]$ and subgroups of $G$ containing $[G,G]$, we get two distinct index-$a_{1}$ subgroups such that the quotients are abelian.
@@
Let $n,r\in\mathbb{N}$ and $p$ be an odd prime. For a prime power $q=p^{r}$, if $q\equiv 1$ $mod$ $n$, then ${\rm GL}_{2}(\mathbb{F}_{q})$ has a unique index-$n$ normal subgroup such that the quotient is abelian. (The quotient with that subgroup is in fact cyclic). This unique subgroup is given by
@@
Proof.
Let $G={\rm GL}_{2}(\mathbb{F}_{q})$. Then $[G,G]=SL_{2}(\mathbb{F}_{q})$ and $G/[G,G]\cong\mathbb{F}_{q}^{\times}$ (isomorphism being through the surjective determinant map) which is cyclic of order $m=q-1$. Now since $q\equiv 1$ $mod$ $n$, we have $n\mid m$. Now consider the determinant map $det:G\rightarrow\mathbb{F}_{q}^{\times}$ with kernel $SL_{2}(\mathbb{F}_{q})$. By applying similar argument to $det$ map as applied to $\pi$ map in previous proposition, we are done.
Remark 2.1.1.
Above corollary holds for ${\rm GL}_{k}(\mathbb{F}_{q})$ as well for any $k\in\mathbb{N}$.
@@
Corollary 2.1.3.
Let $r\in\mathbb{N}$ and $p$ be an odd prime. For a prime power $q=p^{r}$, $SL_{2}(\mathbb{F}_{q})$ is the unique index-$(q-1)$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
@@
    If there exists $x\in G$ such that $x^{n}=1$ and $\{1,x,x^{2},...,x^{n-1}\}$ is a set of representatives for $H$-cosets in $G$, then
    $(n,m/n)=gcd(n,m/n)=1.$
@@
    $1\rightarrow[G,G]\rightarrow G\rightarrow G/[G,G]\rightarrow 1$
@@
    Let $n=m$ (hence $[G,G]$ is the unique index-$m$ subgroup of $G$ such that quotient is abelian). Then the above exact sequence is right split if and only if there exists $x\in G$ such that $x^{m}=1$ and $\{1,x,x^{2},...,x^{m-1}\}$ is a set of representatives for $[G,G]$-cosets in $G$.
@@
    Suppose we have a set of representatives for $H$-cosets in $G$ of the form $\{1,x,x^{2},...,x^{n-1}\}$ with $x\in G$ such that $x^{n}=1$. Let $y$ be a generator of cyclic group $G/[G,G]$ and let $0\leq l\leq(m-1)$ be such that $\pi(x)=y^{l}$. As $x^{n}=1$, we have $1=\pi(x)^{n}=y^{ln}$ and hence $m\mid ln$. Thus we have
    $j=ln/m\in\mathbb{Z}\ \text{and}\ 1\leq j\leq(n-1).$
    Now assume on the contrary that $(n,m/n)\neq 1$. Then for $k=n/(n,m/n)\leq(n-1)$, we have $n\mid km/n$, hence $n\mid kl$, and therefore $\pi(x^{k})=y^{lk}$ is an $n$-th power in $G/[G,G]$, hence $x^{k}\in H$. Hence we get a contradiction. Furthermore, if $(j,n)\neq 1$ then for $k=n/(j,n)$ we have $n\mid lk$, which gives a contradiction. Hence $(j,n)=1$.
@@
    If $(n,m/n)=1$, then we clearly have that $n\nmid km/n$ for all $1\leq k\leq(n-1)$. Consider the quotient map $\pi:G\rightarrow G/[G,G]$. Since given exact sequence is right split, we have an injective homomorphism $\iota:G/[G,G]\rightarrow G$ such that $\pi\circ\iota=id_{G/[G,G]}$. Let $x=\iota(y^{jm/n})\in G$ where $1\leq j\leq(n-1)$ and $(j,n)=1$. Hence $\pi(x)=y^{jm/n}$. Since $\iota$ is injective, $x$ will be an element of order $n$.  Now since $n\nmid km/n$ for all $1\leq k\leq(n-1)$ and since $(j,n)=1$, we have $n\nmid jkm/n$ for all $1\leq k\leq(n-1)$. Hence for all $1\leq k\leq(n-1)$, $\pi(x^{k})=y^{jkm/n}$ is not an $n$-th power in $G/[G,G]$. Hence $x^{k}\not\in H$. Therefore we get a set of representatives for $H$-cosets in $G$ of the form $\{1,x,x^{2},...,x^{n-1}\}$ with $x\in G$ such that $x^{n}=1$.
@@
    Let us assume that the exact sequence is right split. Since $n=m$, we have $(n,m/n)=(m,1)=1$. Hence from (2), we get the required set of representatives for $[G,G]$-cosets in $G$.
    Conversely, if there exists a set of representatives for $[G,G]$-cosets in $G$ of the form $\{1,x,x^{2},...,x^{m-1}\}$ with $x\in G$ such that $x^{m}=1$. Then clearly $\pi(x)=x[G,G]$ is a generator of cyclic group $G/[G,G]$. Define $\iota:G/[G,G]\rightarrow G$ by
    $\iota(\pi(g)):=x^{r}\ \text{if}\ \pi(g)\in G/[G,G]\ \text{is of the form}\ \pi(x)^{r},r\geq 0.$
    Hence $\iota$ is an injective homomorphism satisfying $\pi\circ\iota=id_{G/[G,G]}$ and consequently the exact sequence is right split.
@@
Let $n,r\in\mathbb{N}$. Consider a prime power $q=p^{r}$ such that $q\equiv 1$ mod $n$. Let $H$ be the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian. Then the following are equivalent.
  • $(n,(q-1)/n)=1$.
@@
Let $G={\rm GL}_{2}(\mathbb{F}_{q})$ then $[G,G]=SL_{2}(\mathbb{F}_{q})$ and $G/[G,G]\cong\mathbb{F}_{q}^{\times}$ (isomorphism is defined through the determinant map) is cyclic of order $m=q-1$. Since $q\equiv 1$ $mod$ $n$, we have $n\mid m$. Consider the exact sequence given by the determinant map
$1\rightarrow[G,G]\rightarrow G\xrightarrow{det}\mathbb{F}_{q}^{\times}\rightarrow 1$
Let $\iota:\mathbb{F}_{q}^{\times}\rightarrow G$ be the map defined by
$\iota(a)=\begin{pmatrix}1&0\\ 0&a\end{pmatrix}\ \text{ for all}\ a\in\mathbb{F}_{q}^{\times}.$
This is a homomorphism satisfying $det\circ\iota=id_{\mathbb{F}_{q}^{\times}}$. Hence we conclude that the exact sequence is right split.
By applying similar argument to $det$ map as applied to $\pi$ map in previous proposition, we are done.
Remark 2.2.3.
Above corollary holds for ${\rm GL}_{k}(\mathbb{F}_{q})$ as well for any $k\in\mathbb{N}$. This can be proved using the map
$\iota(a)=\begin{pmatrix}I_{k-1\times k-1}&0\\ 0&a\end{pmatrix}_{k\times k}\ \text{ for all}\ a\in\mathbb{F}_{q}^{\times}.$
@@
$\begin{split}p^{r}\equiv 1\ &mod\ n,\\ (n,(p^{r}-1)/n)&=1.\end{split}$
Proof.
Suppose a prime $p$ satisfies the given conditions. Let $q=p^{r}$. Choose $0\leq k\leq(n-1)$ such that
$\frac{q-1}{n}\equiv k\ mod\ n.$
Observe that $(k,n)=1$ since $(n,(q-1)/n)=1$. Hence $q\equiv kn+1$ $mod$ $n^{2}$.
Conversely, each $p$ satisfying $p^{r}\equiv kn+1$ $mod$ $n^{2}$ for some $0\leq k\leq(n-1)$ with $(k,n)=1$, satisfies given conditions.
Now for each $0\leq k\leq(n-1)$ with $(k,n)=1$, we have $(kn+1,n^{2})=1$. Hence by Dirichlet’s theorem on primes in arithmetic progression, we conclude that for each such $k$, there are infinitely many primes such that
$\begin{split}p&\equiv kn+1\ mod\ n^{2},\\ \text{and hence}\ \ p^{r}&\equiv(kn+1)^{r}\ mod\ n^{2}.\end{split}$
Now for each such $k$, $(rk,n)=1$, because $(k,n)=(r,n)=1$. Hence there are infinitely many primes satisfying given conditions. ∎
@@ 2.2. Galois correspondence, right splitting and Galois theoretic results
For base case $k=2$, see the following diagram LABEL:diagram-compositum.
@@
{tikzpicture}
Let ${\rm Gal}(K/\mathbb{Q})=\{x_{i}H|1\leq i\leq n\}$ where $x_{i}H_{1}=\theta(y_{i}H_{2})$ for all $i$. Here $\{x_{i}\in G_{1}|1\leq i\leq n\}$ and $\{y_{i}\in G_{2}|1\leq i\leq n\}$ are sets of representatives of $H_{1}$ and $H_{2}$ cosets in $G_{1}$ and $G_{2}$ respectively, which are closed under multiplication.
By Galois Correspondence Theorems 2.1 and 2.6 in [2], we have ${\rm Gal}(E_{1}E_{2}/E_{1})\cong H_{2}$, ${\rm Gal}(E_{1}E_{2}/E_{2})\cong H_{1}$ and ${\rm Gal}(E_{1}E_{2}/K)\cong H_{1}\times H_{2}$.
@@
Consider map $\sigma:M\rightarrow G_{1}\times G_{2}$ given by $\sigma(g)=(g_{1},g_{2})$, where $g_{i}=g|_{E_{i}}$ for $i=1,2$. The map $\sigma$ is clearly a well defined injective group homomorphism (see Thm 1.1 [2]).
@@
[…]
where $i(h_{1},h_{2})=(x_{1}h_{1},y_{1}h_{2})$ ($x_{1}\in H_{1},y_{1}\in H_{2}$) and $\pi(x_{i}h_{1},y_{i}h_{2})=x_{i}H_{1}$. The sequence is clearly exact.
@@
Now assume that the statement is true for $k=m$, where $12\leq m\leq l-$ 1, that is,
@@
    In particular, A set of representatives $\{1,x,x^{2},...,x^{n-1}\}$ for $H$-cosets in $G$ which forms a cyclic subgroup of $G$, satisfies the condition. We get a criterion for existence of such a set of representatives in Prop 2.2.
@@
    $1\rightarrow H_{i}\rightarrow G_{i}\rightarrow G_{i}/H_{i}\rightarrow 1$
@@
In the above theorem, we could have assumed a symmetric but stronger condition $E_{1}...E_{i-1}E_{i+1}...E_{n}\cap E_{i}=K$ for all $1\leq i\leq n$ which implies the condition that we have assumed $E_{1}...E_{k}\cap E_{k+1}=K$ for all $11\leq k\leq n-$ 1 since $E_{1}...E_{k}\cap E_{k+1}\subset E_{1}...E_{k}E_{k+2}...E_{n}\cap E_{k+1}=K$. But it would have made our theorem weaker. (Note: Condition for $i=1$ is not required, it is written for symmetry).
@@
    Under the conditions of above theorem, we have $(H_{1}\times H_{2}\times\dots\times H_{l})\rtimes G_{1}/H_{1}\xhookrightarrow{}G_{1}\times G_{2}\times\dots\times G_{l}$ from Thm 1.1 [2].
@@
We now describe an independently interesting consequence of the Galois Correspondence (see Theorems 2.1 and 2.6 in [2].)
@@
Let finite groups $H_{1},H_{2},...,H_{n}$ be Galois groups over $K$ of extensions $E_{1},E_{2},...,E_{n}$ respectively which are contained in $\bar{\mathbb{Q}}$. Fix $k\leq n-1$ and suppose $E_{i_{1}}\cdots E_{i_{j}}\cap E_{i_{j+1}}=K$ for all $11\leq j\leq k-$ 1 where $i_{l}$ are distinct elements from $1$ to $k$. Then the following statements are equivalent.
@@
    $E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{k+1}\cap E_{1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}=E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}$ for any $11\leq i<j\leq k+$ 1.
@@
Now since ${\rm Gal}(D_{i}/D_{ij})\cong H_{j}$ and ${\rm Gal}(D_{j}/D_{ij})\cong H_{i}$, we have by Thm 1.1 [2],

### v2

@@ 1. Introduction
In this article, we establish some instances of Inverse Galois problem. In Sec. 2, through Galois correspondence and right splitting of some exact sequences of groups, we obtain the main results Theorem 2.4, Theorem 2.6. We then apply these general results to the case in [1] and obtain interesting consequences as corollaries which are as follows.
Theorem 1.1.
Let $p$ be a prime $\geq 5$.
  1. (1)
    The group $({\rm SL}_{2}(\mathbb{F}_{p})\times{\rm SL}_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/(p-1)\mathbb{Z}$ (with semidirect product group law as in Thm 2.6) is realizable as Galois group over $\mathbb{Q}$.
  2. (2)
    The group $({\rm PSL}_{2}(\mathbb{F}_{p})\times{\rm PSL}_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/2\mathbb{Z}$ (with semidirect product group law as in Thm 2.4) is realizable as Galois group over $\mathbb{Q}$.
  3. (3)
    For the case when $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ (hence normal) subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{p})/H$ (with semidirect product group law as in Thm 2.6) is realizable as Galois group over $\mathbb{Q}$.
In Sec. 3, using the algebraic operation induction on Galois representations and right splitting of some exact sequences of groups, we obtain Thm. 3.7, Thm. 3.8 as the main results. We then apply these general results to the case in [8] and obtain following interesting consequences in Thm. 3.12 and Thm. 3.13.
Theorem 1.2.
  1. (1)
    For a prime $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realizable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Thm 3.7). The group obtained here is not isomorphic to ${\rm PGL}_{2}(\mathbb{F}_{p})$.
  2. (2)
    The group ${\rm PSL}_{2}(\mathbb{F}_{q})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7) for (1) $q=p$ for $p\geq 5$, (2) $q=p^{2}$ for $p\equiv\pm 2\ mod\ 5,\ p\geq 7$, (3) $q=p^{2}$ for $p\equiv\pm 3\ mod\ 8,\ p\geq 5$, (4) $q=p^{3}$ for odd prime $p\equiv\pm 2,\pm 3,\pm 4,\pm 6\ mod\ 13$ and (5) $q=5^{3},3^{5},3^{4}$.
Then for the algebraic operations direct sum and tensor products on Galois representations we obtain an unanticipated and interesting result Prop 3.15.
Then by using the algebraic operations direct sum and tensor products on Galois representations and right splitting of some exact sequences of groups, we obtain Thm. 3.19 as the main result. We then apply this general result to the case in [8] and obtain that ${\rm PSL}_{2}(\mathbb{F}_{p})\times\mathbb{Z}/2\mathbb{Z}$ is realizable as Galois Group over $\mathbb{Q}$ for $p\geq 5$.
Right Splitting of exact sequences of groups is the common thread that runs through the whole article.
@@ 2.1. Right splitting and group theoretic results
Recall that if $H$ is a normal subgroup of $G$ such that $G/H$ is abelian then $G,G]\subset H$. Following is a simple lemma where the converse follows from structure theorem for abelian groups.
Lemma 2.1.
@@
where $\pi$ is the quotient homomorphism $G\rightarrow G/[G,G]$.
Let $\mathbb{F}_{q}$ be finite field of $q$ elements where $q=p^{r}$ for an odd prime $p$ and $r\in\mathbb{N}$. We know that ${\rm GL}_{2}(\mathbb{F}_{q}),{\rm GL}_{2}(\mathbb{F}_{q})]=SL_{2}(\mathbb{F}_{q})$ and ${\rm GL}_{2}(\mathbb{F}_{q})/SL_{2}(\mathbb{F}_{q})\cong\mathbb{F}_{q}^{\times}$ through the surjective determinant map and $\mathbb{F}_{q}^{\times}$ is cyclic. By applying similar argument to detdet map as applied to $\pi$ map in above lemma, we get the following.
@@
Let $n\in\mathbb{N}$. If $q\equiv 1$ modmod $n$, then ${\rm GL}_{2}(\mathbb{F}_{q})$ has a unique index-$n$ normal subgroup such that the quotient is abelian. (The quotient with that subgroup is in fact cyclic). This unique subgroup is given by
@@
  2. (2)
    $SL_{2}(\mathbb{F}_{q})$ is the unique index-(q1)(q-1) subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
For $a,b\in\mathbb{N}$, we denote $gcd(a,b)$ by $(a,b)$.
@@
    If there exists $x\in G$ such that $x^{n}=1$ and $\{1,x,x^{2},...,x^{n-1}\}$ is a set of representatives for $H$-cosets in $G$, then (n,m/n) =1.
@@
    (2.1) $1\rightarrow[G,G]\rightarrow G\rightarrow G/[G,G]\rightarrow 1$
@@
    Let $n=m$ (hence $[G,G]$ is the unique index-$m$ subgroup of $G$ such that quotient is abelian). Then the exact sequence 2.1 is right split if and only if there exists $x\in G$ such that $x^{m}=1$ and $\{1,x,x^{2},...,x^{m-1}\}$ is a set of representatives for $[G,G]$-cosets in $G$.
@@
    Suppose we have a set of representatives for $H$-cosets in $G$ of the form $\{1,x,x^{2},...,x^{n-1}\}$ with $x\in G$ such that $x^{n}=1$. Let $y$ be a generator of cyclic group $G/[G,G]$ and let $0\leq l\leq(m-1)$ be such that $\pi(x)=y^{l}$. As $x^{n}=1$, we have $1=\pi(x)^{n}=y^{ln}$ and hence $m\mid ln$. Thus we have j = ln/m ∈Z and 1≤j≤(n-1).
    Now assume on the contrary that $(n,m/n)\neq 1$. Then for $k=n/(n,m/n)\leq(n-1)$, we have $n\mid km/n$, hence $n\mid kl$, and therefore $\pi(x^{k})=y^{lk}$ is an $n$-th power in $G/[G,G]$, hence $x^{k}\in H$. Hence we get a contradiction. Similarly we can show that $(j,n)=1$.
@@
    If $(n,m/n)=1$, then we clearly have that $n\nmid km/n$ for all $1\leq k\leq(n-1)$. Consider the quotient map $\pi:G\rightarrow G/[G,G]$. Since given exact sequence 2.1 is right split, we have an injective homomorphism $\iota:G/[G,G]\rightarrow G$ such that $\pi\circ\iota=id_{G/[G,G]}$. Let $x=\iota(y^{jm/n})\in G$ where $1\leq j\leq(n-1)$ and $(j,n)=1$. Hence $\pi(x)=y^{jm/n}$. Since $\iota$ is injective, $x$ will be an element of order $n$. Now since $n\nmid km/n$ for all $1\leq k\leq(n-1)$ and since $(j,n)=1$, we have $n\nmid jkm/n$ for all $1\leq k\leq(n-1)$. Hence for all $1\leq k\leq(n-1)$, $\pi(x^{k})=y^{jkm/n}$ is not an $n$-th power in $G/[G,G]$. Hence $x^{k}\not\in H$. Therefore we get the required set of representatives for $H$-cosets in $G$.
@@
    If the exact sequence 2.1 is right split, then by (2), we are done. The converse is easy.
@@
Let $n\in\mathbb{N}$. Consider $q$ such that $q\equiv 1$ mod $n$. Let $H$ be the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian. Then the following are equivalent.
  • (n,(q1)/n)=1(n,(q-1)/n)=1.
@@
Let $G={\rm GL}_{2}(\mathbb{F}_{q})$. We know that below exact sequence given by the determinant map is right split.
(2.2) $1\rightarrow[G,G]\rightarrow G\xrightarrow{det}\mathbb{F}_{q}^{\times}\rightarrow 1.$
By applying similar argument to detdet map as applied to $\pi$ map in previous proposition, we are done.∎
The following follows from Dirichlet’s theorem on primes in arithmetic progression.
@@
$p^{r}\equiv 1\ mod\ n\ \text{and}\ (n,(p^{r}-1)/n)=1.$
@@ 2.2. Galois correspondence, right splitting and Galois theoretic results
For base case $k=2$, see the following diagram.
@@
$\mathbb{Q}$$E_{1}\cap E_{2}$$E_{2}$$E_{1}E_{2}$$E_{1}$$H_{2}$$H_{1}$$H_{2}$$H_{1}$$H_{1}\times H_{2}$$G_{1}/H_{1}$$G_{1}$$G_{2}$$M$
Let ${\rm Gal}(K/\mathbb{Q})=\{x_{i}H_{1}|1\leq i\leq n\}$ where $x_{i}H_{1}=\theta(y_{i}H_{2})$ for all $i$. Here $\{x_{i}\in G_{1}|1\leq i\leq n\}$ and $\{y_{i}\in G_{2}|1\leq i\leq n\}$ are sets of representatives of $H_{1}$ and $H_{2}$ cosets in $G_{1}$ and $G_{2}$ respectively, which are closed under multiplication.
By Galois Correspondence [2, Theorem 2.1] and [2, Theorem 2.6], we have ${\rm Gal}(E_{1}E_{2}/E_{1})\cong H_{2}$, ${\rm Gal}(E_{1}E_{2}/E_{2})\cong H_{1}$ and ${\rm Gal}(E_{1}E_{2}/K)\cong H_{1}\times H_{2}$.
@@
Consider map $\sigma:M\rightarrow G_{1}\times G_{2}$ given by $\sigma(g)=(g_{1},g_{2})$, where $g_{i}=g|_{E_{i}}$ for $i=1,2$. The map $\sigma$ is clearly a well defined injective group homomorphism (see [2, Theorem 1.1]).
@@
(2.3) $1\rightarrow(H_{1}\times H_{2})\xrightarrow{i}\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})\xrightarrow{\pi}G_{1}/H_{1}\rightarrow 1$
where $i(h_{1},h_{2})=(x_{1}h_{1},y_{1}h_{2})$ ($x_{1}\in H_{1},y_{1}\in H_{2}$) and $\pi(x_{i}h_{1},y_{i}h_{2})=x_{i}H_{1}$. Because of our multiplicatively closed assumption, $x_{1}$ and $y_{1}$ are identities of $G_{1}$ and $G_{2}$ respectively. Hence $i$ is injective. Also $\pi$ is surjective and $ker(\pi)=Image(i)$. Thus the sequence is exact.
@@
Now assume that the statement is true for $k=m$, where $2\leq m\leq l-1$, that is,
@@
    In particular, A set of representatives $\{1,x,x^{2},...,x^{n-1}\}$ for $H$-cosets in $G$ which forms a cyclic subgroup of $G$, satisfies the condition. We get a criterion for existence of such a set of representatives in Proposition 2.2.
@@
    (2.4) $1\rightarrow H_{i}\rightarrow G_{i}\rightarrow G_{i}/H_{i}\rightarrow 1$
@@
In the above theorem, we could have assumed a symmetric but stronger condition
$E_{1}...E_{i-1}E_{i+1}...E_{n}\cap E_{i}=K$ for all $1\leq i\leq n$ which implies the condition that we have assumed $E_{1}...E_{k}\cap E_{k+1}=K$ for all $1\leq k\leq n-1$ since $E_{1}...E_{k}\cap E_{k+1}\subset E_{1}...E_{k}E_{k+2}...E_{n}\cap E_{k+1}=K$. But it would have made our theorem weaker. (Note: Condition for $i=1$ is not required, it is written for symmetry).
@@
    Under the conditions of above theorem, we have $(H_{1}\times H_{2}\times\dots\times H_{l})\rtimes G_{1}/H_{1}\xhookrightarrow{}G_{1}\times G_{2}\times\dots\times G_{l}$ from [2, Theorem 1.1].
@@
We now describe an independently interesting consequence of the Galois Correspondence (see [2, Theorem 2.1] and [2, Theorem 2.6].)
@@
Let finite groups $H_{1},H_{2},...,H_{n}$ be Galois groups over $K$ of extensions $E_{1},E_{2},...,E_{n}$ respectively which are contained in $\bar{\mathbb{Q}}$. Fix $k\leq n-1$ and suppose $E_{i_{1}}\cdots E_{i_{j}}\cap E_{i_{j+1}}=K$ for all $1\leq j\leq k-1$ where $i_{l}$ are distinct elements from $1$ to $k$. Then the following statements are equivalent.
@@
    $E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{k+1}\cap E_{1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}=E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}$ for any $1\leq i<j\leq k+1$.
@@
Now since ${\rm Gal}(D_{i}/D_{ij})\cong H_{j}$ and ${\rm Gal}(D_{j}/D_{ij})\cong H_{i}$, we have by [2, Theorem 1.1],